# An analysis of nuclear parton distribution function based on Kullback-Leibler divergence

### v1

@@ IV Framework and results
The KL divergence can be viewed as a measure of the inefficiency of assuming that the distribution is $q(x)$ when the true distribution is $p(x)$. Here, we effectively use the deuteron’s structure function to describe that of nucleus A. Qualitatively speaking, as A increases, nuclear modification becomes more apparent, leading to an increase in the KL divergence — a trend that is consistent with the results shown in Table 1.
These results essentially quantify the findings which have already been known. More importantly, we are interested in whether the KL divergence can provide new insights into nucleon structure, especially in the EMC region. Prior to the discovery of the EMC effect, it was commonly believed that the ratio $(F_{2}^{A}(x,Q^{2})/A)/(F_{2}^{D}(x,Q^{2})/2)$ would be unity in the intermediate-$x$ region, implying a corresponding KL divergence of zero. However, the EMC effect revealed that this ratio exhibits a deviation from unity, showing a monotonic decrease in the approximate $x$-range of 0.25 to 0.65.
Now, for a given nucleus A, if we fix the two endpoints ($x=0.25$ and $0.65$) respectively. The question is: can we establish a certain criterion to determine the shape of the structure function within the EMC region? This seems difficult since we are trying to construct this function with insufficient data — neither perturbative QCD nor lattice QCD can be applied on this issue. Interestingly, the problem encountered here bears a striking resemblance to the famous brachistochrone problem proposed by Bernoulli in 1696 brachistochronecurve . Both involve finding the extremum of a functional under fixed boundary conditions — the brachistochrone problem seeks the minimal time of descent, while our task requires establishing an extremum principle for a physical observable.
We assume that, with the endpoints fixed, the actual structure function of the nucleus tends to minimize the KL divergence in Eq. (9), which we refer to as the “minimum relative entropy” hypothesis. The $\bar{p}^{A}(x)$ represents the structure function we aim to determine. First, we adopt a polynomial form of parametrization
Using the EPPS21 global fits, we can derive constraints at the endpoints, which determine two parameters in Eq. (10), while the remaining two are fixed by the minimum relative entropy hypothesis. In Figure 1, we provide two-dimensional visualization of the KL divergence as a function of $a_{1}$ and $b_{1}$ parameters. The KL divergence forms a distribution in the parameter space that allows us to pin down its minimum point. The position marked with a red dot indicates the corresponding $a_{1}$ and $b_{1}$ with minimum KL divergence.
Figure 1: The KL divergence as a function of $a_{1}$ and $b_{1}$. The red dot indicates the position with minimum KL divergence for 4He, 12C, 56Fe and 208Pb, at the momentum transfer $Q^{2}=10\,{\rm GeV}^{2}$. The orange and magenta dots are two selected reference points which will be useful later.
Appendix A collects the four parameters ($a_{1}$, $b_{1}$, $c_{1}$ and $d_{1}$) in Eq. (10), which are determined by minimizing the KL divergences in Eq. (9) for different nuclei. The second column of Table 2 collects the calculated minimum KL divergence from Eq. (IV).
Table 2: The calculated minimum KL divergences for different nuclei with the polynomial and canonical parameterizations in intermediate-$x$ region $[0.25,0.65]$.
   Nucleus   $D_{\mathrm{KL}}^{\textrm{EMC}}(\bar{p}^{A}_{\textrm{poly}}\|q^{d})$   $D_{\mathrm{KL}}^{\textrm{EMC}}(\bar{p}^{A}_{\textrm{can}}\|q^{d})$
3He $1.3\times 10^{-5}$ $1.2\times 10^{-5}$
4He $3.0\times 10^{-5}$ $2.9\times 10^{-5}$
9Be $7.0\times 10^{-5}$ $6.9\times 10^{-5}$
12C $5.0\times 10^{-5}$ $5.0\times 10^{-5}$
56Fe $1.21\times 10^{-4}$ $1.21\times 10^{-4}$
197Au $2.52\times 10^{-4}$ $2.58\times 10^{-4}$
208Pb $2.63\times 10^{-4}$ $2.70\times 10^{-4}$
Figure 2: The obtained structure function $\bar{p}^{A}_{\textrm{poly}}(x)$ with minimum KL divergence (red line), the structure function $p^{A}(x)$ given by the global fit of experimental data (black line), and the two reference structure functions $\bar{p}^{A}_{\textrm{ref1}}(x)$ and $\bar{p}^{A}_{\textrm{ref2}}(x)$ (orange and magenta lines).
To quantify the perceived separation between various structure functions and the QCD global results in Figure 2, we make use of the $L^{2}$ norm which is defined as
here $p^{A}(x)$ is the structure function from global fit. The results are listed in the second column of Table 3. It is evident that the norm calculated from $\bar{p}^{A}_{\textrm{poly}}(x)$ is much less than the reference curves $p^{A}_{\textrm{ref1}}(x)$ and $p^{A}_{\textrm{ref2}}(x)$, indicating the minimum relative entropy hypothesis might be a reliable criterion in constructing nPDFs of bound nucleon from PDFs of free nucleon. It is worth noting that the norms $N(\bar{p}^{A}_{\textrm{poly}})$ generally exhibit an increasing trend with the increasing of mass number $A$.
3He $6.61\times 10^{-4}$ $2.11\times 10^{-2}$ $2.33\times 10^{-2}$
4He $1.79\times 10^{-3}$ $50\times 10^{-2} 2.2$$25\times 10^{-2}$
9Be $2.38\times 10^{-3}$ $4.07\times 10^{-2}$ $2.91\times 10^{-2}$
12C $2.20\times 10^{-3}$ $49\times 10^{-2} 2.2$$26\times 10^{-2}$
56Fe $3.10\times 10^{-3}$ $1.22\times 10^{-2}$ $3.03\times 10^{-2}$
197Au $4.11\times 10^{-3}$ $1.34\times 10^{-2}$ $2.31\times 10^{-2}$
208Pb $4.14\times 10^{-3}$ $2.53\times 10^{-2}$ $2.89\times 10^{-2}$
There is an issue regarding the dependence on the functional form of the parametrization. A reasonable test is necessary. In Eq. (10), we adopted a polynomial form and expanded it to the cubic term $x^{3}$. We have also extended the expansion to the quartic term $x^{4}$, the resulting $\bar{p}^{A}_{\textrm{poly}}(x)$ remained identical to the cubic case. Moreover, the widely used canonical parametrization form with exponential and power-law endpoint behaviour were also tested,
@@
Appendix A collects the four parameters ($a_{2}$, $b_{2}$, $c_{2}$ and $d_{2}$) in Eq. (13), which are determined by minimizing the KL divergences for different nuclei.
The third column of Table 2 collects the calculated minimum KL divergence from Eq. (IV). As one can see, the obtained values show remarkable consistency with the minimum KL divergence calculated through Eq. (IV), consequently yielding nearly identical norms in Table 3. This agreement becomes more visually discernible when displaying the structure functions together. Figure 3 shows the two kinds of structure functions obtained from the polynomial and canonical form of parameterizations, as well as the global fitting results from EPPS21 set. Visually indistinguishable differences are observed between the structure functions obtained from these two parameterizations. These results demonstrate the robustness of the calculated minimum KL divergence across different parametrization forms. When the two endpoints in the intermediate-$x$ region are fixed, both the value of the minimum KL divergence and the shape of the nuclear structure function are almost exclusively determined by the minimum relative entropy hypothesis. It complements the traditional global analysis on nPDFs. We note that for nuclei with large mass number, the structure functions obtained from the KL divergence approach exhibit apparent deviations from those from global fits, suggesting the need for additional study. In the next section, we will investigate the performance of this method when applied to gluon nPDFs.
Figure 3: The obtained structure function $\bar{p}^{A}_{\textrm{poly}}(x)$ from polynomial form of parameterization (red line), the structure function $\bar{p}^{A}_{\textrm{can}}(x)$ from canonical form of parametrization (blue dashed line), as well the structure function $p^{A}(x)$ given by the global fit of experimental data (black line).
V Determining gluon nPDFs through KL divergence
The nucleons are formed by constituent quarks and gluons, the exploration of distributions of mediators of strong interaction — gluons in nuclei would be also important. Furthermore, our understanding of gluon nPDFs remains much poorer compared to quark nPDFs. Whether gluons also exhibit an EMC effect is an open question under active investigation Chen:2016bde ; Wang:2016mzo ; Xu:2019wso ; Hatta:2019ocp ; Wang:2020uhj ; Wang:2021elw ; Wang:2024ikx ; Yang:2023zmr ; Wang:2022kwg ; Huang:2021cac ; Wang:2024cpx ; Hu:2021fxa . Our results in the last section show that the minimum relative entropy hypothesis gives the reasonable quark nPDFs. We now apply this approach to determine the gluon nPFDs in the intermediate-$x$ region. The KL divergence between the gluon nPDFs of nucleus A and deuteron is defined as
$D_{\mathrm{KL},g}^{\textrm{EMC}}(p^{A}_{g}\|q_{g}^{d})\!=\!\int_{0.25}^{0.50}\left.p_{g}^{A}(x)\right|_{\textrm{nor}}\ln\frac{\left.p_{g}^{A}(x)\right|_{\textrm{nor}}}{\left.q_{g}^{d}(x)\right|_{\textrm{nor}}}\,dx.$ (15)
Here $p_{g}^{A}(x)$ is taken as the true distribution, which is the gluon nPDF for nucleus A, we still utilize its results from global analysis given by EPPS21 Collaboration. The endpoints 0.25 and 0.50 are also derived from the EPPS21 set. $q_{g}^{d}(x)$ is taken as the reference distribution, which is the gluon nPDF for deuteron, whose value is taken from the CT18A set. These distributions necessitate normalization analogous to that in Eq. (7)
$\left\{\begin{aligned} &\left.p_{g}^{A}(x)\right|_{\textrm{nor}}=\frac{1}{\int_{0.25}^{0.50}p_{g}^{A}(x)dx}\,p_{g}^{A}(x)=z_{p,g}\,p_{g}^{A}(x)\,,\\ &\left.q_{g}^{d}(x)\right|_{\textrm{nor}}=\frac{1}{\int_{0.25}^{0.50}q_{g}^{d}(x)dx}\,q_{g}^{d}(x)=z_{q,g}\,q_{g}^{d}(x)\,.\end{aligned}\right$. (16)
The second column of Table 4 presents the KL divergences for different nuclei.
Table 4: The KL divergences given by the global analysis, as well as the calculated minimum KL divergences with the polynomial and canonical form of parameterizations in the intermediate-$x$ region $[0.25,0.50]$. The four-momentum transfer is chosen as $Q^{2}=10\,\textrm{GeV}^{2}$.
3He $1.54\times 10^{-5}$ $3.34\times 10^{-6}$ $3.44\times 10^{-6}$
4He $2.31\times 10^{-4}$ $4.66\times 10^{-5}$ $4.86\times 10^{-5}$
9Be $3.24\times 10^{-4}$ $6.51\times 10^{-5}$ $7.19\times 10^{-5}$
12C $3.65\times 10^{-4}$ $33\times 10^{-5} 7.7$$70\times 10^{-5}$
56Fe $6.91\times 10^{-4}$ $1.37\times 10^{-4}$ $1.98\times 10^{-4}$
197Au $1.17\times 10^{-3}$ $2.28\times 10^{-4}$ $4.21\times 10^{-4}$
208Pb $1.19\times 10^{-3}$ $2.33\times 10^{-4}$ $3.95\times 10^{-4}$
With the gluon EMC region endpoints fixed at $x=0.25$ and $0.50$, we determine the shape of the gluon nPDFs $\bar{p}_{g}^{A}(x)$ that minimizes the relative entropy
$D_{\mathrm{KL},g}^{\textrm{EMC}}(\bar{p}_{g}^{A}\|q_{g}^{d})=\int_{0.25}^{0.50}\!\!\left.\bar{p}_{g}^{A}(x)\right|_{\textrm{nor}}\ln\frac{\left.\bar{p}_{g}^{A}(x)\right|_{\textrm{nor}}}{\left.q_{g}^{d}(x)\right|_{\textrm{nor}}}\,dx.$ (17)
Simlilar to the quark case, the polynomial and canonical form of parameterizations are utilized to calculate $\bar{p}_{g}^{A}(x)$
In Figure 4, we provide two-dimensional visualization of the KL divergence as a function of $a_{1}^{g}$ and $b_{1}^{g}$ parameters in the polynomial parametrization. In this figure, the position marked with a red dot indicates the corresponding $a_{1}^{g}$ and $b_{1}^{g}$ with minimum KL divergence. Here, visualization of the canonical parametrization is omitted for conciseness.
Next, we present the explicit results of 4He, 12C, 56Fe, 208Pb. For the polynomial parametrization:
and
$\bar{p}^{{}^{208}\textrm{Pb}}_{\textrm{can}}(x)$ $\!=\!$ $\big{(}0.002\,x^{-2.36}(1-x)^{10.92}e^{23.50x}\big{)}\,q^{d}(x).$
for the canonical parametrization. Appendix A collects the calculated parameters ($a_{1}^{g}$, $b_{1}^{g}$, $c_{1}^{g}$ and $d_{1}^{g}$) in Eq. (18) and ($a_{2}^{g}$, $b_{2}^{g}$, $c_{2}^{g}$ and $d_{2}^{g}$) in Eq. (19).
The $L^{2}$ norm is utilized to quantify the differences between parametrization results and the QCD global results,
$N(\bar{p}^{A}_{g,\textrm{poly}})=\sqrt{\int_{0.25}^{0.50}\Big{(}\bar{p}^{A}_{g,\textrm{poly}}(x)-p_{g}^{A}(x)\Big{)}^{2}dx}\,$, (22)
The value of norms are collected in Table 5. It can be seen that the norms given by the two parameterizations differ very little. This supports the independence of the results with respect to the form of parameterization.
Table 5: The norms of structure functions with minimum KL divergences in polynomial (the second column) and canonical (the third column) parameterizations.
   Nucleus   $N(\bar{p}^{A}_{g,\textrm{poly}})$   $N(\bar{p}^{A}_{g,\textrm{can}})$
3He $1.27\times 10^{-3}$ $1.25\times 10^{-3}$
4He $4.70\times 10^{-3}$ $4.58\times 10^{-3}$
9Be $5.54\times 10^{-3}$ $5.45\times 10^{-3}$
12C $5.87\times 10^{-3}$ $5.70\times 10^{-3}$
56Fe $8.02\times 10^{-3}$ $9.05\times 10^{-3}$
197Au $03\times 10^{-2} 1.1$$15\times 10^{-2}$
208Pb $1.05\times 10^{-2}$ $1.28\times 10^{-2}$
Figure 4: The KL divergence as a function of $a_{1}^{g}$ and $b_{1}^{g}$. The red dot indicates the position with minimum KL divergence for 4He, 12C, 56Fe and 208Pb, at the momentum transfer $Q^{2}=10\,{\rm GeV}^{2}$.
Figure 5: The obtained structure function $\bar{p}^{A}_{g,\textrm{poly}}(x)$ with minimum KL divergence in polynomial parametrization (red line), the structure function $\bar{p}^{A}_{g,\textrm{can}}(x)$ with minimum KL divergence in canonical parametrization (blue dashed line), as well as the structure function $p_{g}^{A}(x)$ given by the global fit of experimental data (black line).
With the results in Eq. (V) and (V), we plot the gluon nPDFs given by the minimum relative entropy hypothesis in Figure 5. It can be observed that the gluon nPDF shapes given by the two parameterizations are highly consistent, with errors between them barely discernible. Moreover, both results agree well with the global fits of experimental data, falling within the error bar of EPPS21 result. It is worth noting that the gluon nPDF has relatively large uncertainties due to our currently poor knowledge of it. Therefore, the minimum relative entropy hypothesis may prove more useful in determining the gluon nPDF than in constraining the quark nPDFs.
We find the results in this work intriguing which worth further investigation to uncover the profound physical connections between PDFs and nPDFs within the framework of quantum information theory. The behavior of nonperturbative QCD is very complex. In contrast to conventional approaches, the idea of entropy offers a promising avenue for elucidating the fundamental symmetries and underlying dynamics of QCD that are otherwise hidden. It is remarkable that the minimum relative entropy hypothesis yields a unified and reasonable description of both quark and gluon nPDFs in the intermediate-$x$ region, which will mitigate the artificiality in presupposing nPDF functional forms during global fitting procedures. The present work also has aspects that require further improvement. For instance, the current selection of endpoints in the EMC region relies on experimental results. Additionally, extending the applicable range from the EMC region to the entire region remains unknown — we leave such questions to future study.
VI Concluding remarks
In this work, beyond conventional observables, we introduced the KL divergence, which is a well-known estimators of quantum information theory, as a measure to probe the parton structure of nucleon bound in nucleus. The obtained results are consistent with the latest global fits of experimental data, revealing promising potential for this interdisciplinary research.
In particle physics, determination of nonperturbative inputs apart from the global fit is not only a complementarity but also a necessity to understand the nature of hadrons. The introduction of KL divergence and minimum relative entropy hypothesis offers novel perspective on old problems. It can be used for searching more details of the nuclear structure function and parton distributions. Admittedly, all the results presented in this work should rather be considered primarily as indicative evidence, while the actual potential of KL divergence remains inconclusive. It is hoped that this work could inspire further research on applying information theory to nucleon structure studies.
Acknowledgements
Appendix A Collection of determined parameters
In this appendix, we collect the determined parameters in the polynomial and canonical parameterizations. For the quark case, we give the four parameters ($a_{1}$, $b_{1}$, $c_{1}$ and $d_{1}$) in Eq. (10) and four parameters ($a_{2}$, $b_{2}$, $c_{2}$ and $d_{2}$) in Eq. (13), which are determined by minimizing the KL divergences in Eq. (9) for different nuclei.
@@ Appendix A Collection of determined parameters
Table 6: For the quark case, the determined parameters in polynomial form in Eq. (10) and canonical form in Eq. (13) for different nuclei. The integration interval is $x\in[0.25,0.65]$, as stated in Eq. (9).
@@
For the gluon case, we give the four parameters ($a_{1}^{g}$, $b_{1}^{g}$, $c_{1}^{g}$ and $d_{1}^{g}$) in Eq. (18) and four parameters ($a_{2}^{g}$, $b_{2}^{g}$, $c_{2}^{g}$ and $d_{2}^{g}$) in Eq. (19), which are determined by minimizing the KL divergences in Eq. (17) for different nuclei.
Table 7: For the gluon case, the determined parameters in polynomial form in Eq. (18) and canonical form in Eq. (19) for different nuclei. The integration interval is $x\in[0.25,0.50]$, as stated in Eq. (17).

### v2

@@ Appendix A Different $x$-region KL divergence calculation details
Here, we utilize the nPDF sets from EPPS21 and CT18A. For full-$x$ region $10^{-6},0.99]$ and intermediate-$x$ region $[0.25,0.65]$, we take the $F_{2}^{A}(x,Q^{2})/A$ as $p^{A}(x)$ in Eq. (6) and the $F_{2}^{d}(x,Q^{2})/2$ as $q^{d}(x)$. Qualitatively speaking, as A increases, nuclear modification becomes more apparent, leading to an increase in the KL divergence—a trend that is consistent with the results shown in Table 7. However, it is worth noting that the value for $^{12}\textrm{C}$} is lower than for $^{9}\textrm{Be}$}, a pattern consistent with findings in studies of nucleon short-range correlations which show a non-monotonic dependence on A for $^{12}\textrm{C}$} and $^{9}\textrm{Be}$} CLAS:2019vsb . This may arise from the unique nature of beryllium, whose density is similar to helium, yet whose EMC effect closely resembles that of denser nuclei due to its significant cluster structure.
The values of normalization factors $z_{p}$ and $z_{q}$ defined in Eq. (8) are listed in Table 8.
Table 8: The normalization factors $z_{p}$ and $z_{q}$ for different nuclei in full-$x$ region $10^{-6},0.99]$ and intermediate-$x$ region $[0.25,0.65]$. The four-momentum transfer is chosen as $Q^{2}=10\,\textrm{GeV}^{2}$.
   Nucleus   $z_{p}^{\textrm{full}}$   $z_{p}^{\textrm{EMC}}$
2D($z_{q}$) $7.06$ $19.61$
3He $6.70$ $18.16$
4He $7.12$ $20.32$
9Be $7.26$ $21.05$
12C $7.14$ $20.48$
56Fe $7.27$ $21.16$
197Au $7.48$ $22.20$
208Pb $7.50$ $22.29$
Table 9 presents the normalization factors used in the calculation of the gluon KL divergence.
Table 9: The normalization factors $z_{p,g}$ and $z_{q,g}$ for different nuclei in intermediate-$x$ region $[0.25,0.5]$. The four-momentum transfer is chosen as $Q^{2}=10\,\textrm{GeV}^{2}$.
   Nucleus   $z_{p,g}^{\textrm{EMC}}$
2D($z_{q,g}$) $7.67$
3He $7.74$
4He $7.93$
9Be $7.97$
12C $7.98$
56Fe $8.09$
197Au $8.20$
208Pb $8.21$
Table 10 presents the normalization factors used in the calculation of the nNNPDF30 gluon KL divergence .
Table 10: The normalization factors $z_{p,g}$ and $z_{q,g}$ for different nuclei in intermediate-$x$ region $[0.3,0.55]$. The four-momentum transfer is chosen as $Q^{2}=10\,\textrm{GeV}^{2}$.
   Nucleus   $z_{p,g}^{\textrm{EMC}}$
2D($z_{q,g}$) $13.03$
4He $12.99$
9Be $12.79$
12C $12.65$
56Fe $11.94$
197Au $12.32$
208Pb $12.48$
Appendix B Collection of determined parameters
@@
In this appendix, we collect the determined parameters in the polynomial and canonical parameterizations. For the quark case, we give the four parameters ($a_{1}$, $b_{1}$, $c_{1}$ and $d_{1}$) in Eq. (11) and four parameters ($a_{2}$, $b_{2}$, $c_{2}$ and $d_{2}$) in Eq. (13), which are determined by endpoints constrains and minimizing the KL divergences in Eq. (10) for different nuclei.
@@
Table 11: For the quark case, the determined parameters in polynomial form in Eq. (11) and canonical form in Eq. (13) for different nuclei. The integration interval is $x\in[0.25,0.65]$.
@@
For the gluon case, after applying the EPPS21 global fitting for gluon nPDFs and the minimum relative entropy hypothesis, we give the four parameters ($a_{1}^{g}$, $b_{1}^{g}$, $c_{1}^{g}$ and $d_{1}^{g}$) in Eq. (V) and four parameters ($a_{2}^{g}$, $b_{2}^{g}$, $c_{2}^{g}$ and $d_{2}^{g}$) in Eq. (V), which are determined by endpoints constrains and minimizing the KL divergences in Eq. (18) for different nuclei.
Table 12: For the gluon case, the determined parameters in polynomial form in Eq. (19) and canonical form in Eq. (20) for different nuclei by utilizing the EPPS21 global fits. The integration interval is $x\in[0.25,0.50]$.
@@
For the gluon case, after applying the nNNPDF3.0 global fitting for gluon nPDFs and the minimum relative entropy hypothesis, we give the four parameters ($a_{1}^{g}$, $b_{1}^{g}$, $c_{1}^{g}$ and $d_{1}^{g}$) in Eq. (V) and four parameters ($a_{2}^{g}$, $b_{2}^{g}$, $c_{2}^{g}$ and $d_{2}^{g}$) in Eq. (V), which are determined by endpoints constrains and minimizing the KL divergences in Eq. (18) for different nuclei.
Table 13: For the gluon case, the determined parameters in polynomial form in Eq. (V) and canonical form in Eq. (V) for different nuclei by utilizing the nNNPDF3.0 global fits. The integration interval is $x\in[0.3,0.55]$.
   Nucleus     $a_{1}^{g}$     $b_{1}^{g}$     $c_{1}^{g}$     $d_{1}^{g}$     $a_{2}^{g}$     $b_{2}^{g}$     $c_{2}^{g}$     $d_{2}^{g}$
4He 8.97-8.97 $10.96$ 4.41-4.41 $1.60$ $0.0976$ 1.03-1.03 $2.22$ $6.33$
9Be 32.70-32.70 $39.96$ 16.06-16.06 $3.17$ $0.0060$ 2.24-2.24 $5.44$ $14.74$
12C 47.08-47.08 $57.52$ 23.11-23.11 $4.13$ $0.0014$ 2.86-2.86 $7.16$ $19.15$
56Fe 216.67-216.67 $263.53$ 105.39-105.39 $15.07$ $0.0025$ 2.20-2.20 $15.47$ $30.52$
197Au 359.78-359.78 $435.71$ 173.42-173.42 $23.93$ $1.58*10^{-6}$ 4.96-4.96 $30.91$ $62.49$
208Pb 389.23-389.23 $470.88$ 187.19-187.19 $25.70$ $0.0002$ 2.60-2.60 $31.26$ $56.80$